You have 500 mL of a buffer solution containing 0.3 M Acetic acid and 0.2 M acetate . The PH of the buffer is 4.57. What will be |Delta PH| of this solution after the addition of 20 mL of 1 M KOH? Pka of the weak acid is 4.75
1
Expert's answer
2019-08-26T05:45:41-0400

The additional volume of KOH will react with acetic acid in the buffer according the following equation:

CH3COOH + KOH = CH3COOK + H2O

We have to calculate the amount of added KOH in moles:

"n(KOH)=V(KOH)*C_{KOH}=0.02L*1mol\/L=0.02mol"

We also have to calculate the amounts in moles of CH3COOH and acetate anion in the buffer:

"n(acetate)=C_{acetate}*V_{buffer}=0.2mol\/L*0.5L=0.1mol"

"n(acid)=C_{acid}*V_{buffer}=0.3mol\/L*0.5L=0.15mol"

One can easily note, that some amount of acid will be consumed, while that of acetate anion will be increase:

"n(acetate)_{new}=n(acetate)+n(KOH)=0.10mol+0.02mol=0.12mol"

"n(acid)_{new}=n(acid)-n(KOH)=0.15mol-0.02mol=0.13mol"

In the first approximation, the total volume of the new buffer will be the additive of the old volume and the added one:

"V(buffer)_{new}=V(buffer)+V(KOH)=0.500L+0.020L=0.520L"

Let us now calculate the novel concentrations of acid and acetate anion:

"C(acid)_{new}=\frac{n(acid)_{new}}{V(buffer)_{new}}=\frac{0.13mol}{0.520L}=0.25M"

"C(acetate)_{new}=\frac{n(acetate)_{new}}{V(buffer)_{new}}=\frac{0.12mol}{0.520L}=0.23M"

Now it is right time to evaluate the new pH value:

"pH_{new}=pK_{a}+lg(\frac{C(acetate)_{new}}{C(acid)_{new}})=4.75+lg(\frac{0.23M}{0.25M})=4.71"

Thus, the absolute value of the pH delta is:

|"\Delta" pH| = 4.71 - 4.57 = 0.14